From the following information calculate national income and domestic income.(5)
Items. Amount
I. Wages. 10000
II. Rent. 5000
III. Interest. 400
IV. Dividend. 3000
V. Mixed income. 400
VI. Undistributed profit. 200
VII. Social security contribution. 400
VIII. Corporate profit tax. 400
IX. N.F.I.A. 1000
1
Expert's answer
2020-08-11T08:46:37-0400

Domestic Income = (Wages + Rent + Interest + Dividend + Mixed income + Undistributed profit + Social security contribution + Corporate profit tax)


Domestic income = (10000 + 5000 + 400 + 3000 + 400 + 200 + 400 + 400) = 19800


Domestic income = 19800


National income = Domestic income + N. F. I. A

National income = (19800 + 1000) = 20800


National Income = 20800